Opened 7 years ago

Closed 4 years ago

#451 closed flight processing (fixed)

BGS11/01, flight day 236a/2012, East Greenland

Reported by: knpa Owned by: knpa
Priority: alpha 4 high Milestone: 2012 data processing completion
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by mark1)

Data location: /users/rsg/arsf/arsf_data/2012/flight_data/greenland_iceland/BGS11_01-2012_236a_East_Greenland/

Data arrived from ARSF via SATA disk CAM11 24/09/2012

BGS11/01 Scientific objective: unknown

BGS11/01 Priority: ?

PI: Tapsa Tukiainer

Sensors:

Camera (16/11/2012)
Eagle (16/11/2012)
Hawk (16/11/2012)
LiDAR (04/01/2013)

Change History (37)

comment:1 Changed 7 years ago by knpa

Eagle 1 and 2 raws are currently missing (they were corrupted on the disk received). Have asked for them to be re-sent.

comment:2 Changed 7 years ago by knpa

  • Description modified (diff)
  • Priority changed from alpha 4 medium to alpha 4 high

comment:3 Changed 7 years ago by besm

Starting nav processing

comment:4 Changed 7 years ago by besm

Stopping nav processing - turns out we don't have the raw nav files yet.

Starting RCD processing

comment:5 Changed 7 years ago by mark1

Processed the navigation using PPP - should be good enough to get on with timing correcting the hyperspectral data. When base station data arrives we can reprocess the nav if required.

comment:6 Changed 7 years ago by mark1

Completed the timing correction of hyperspectral data using the PPP data.

comment:7 Changed 7 years ago by knpa

  • Description modified (diff)

comment:8 Changed 7 years ago by anhi

Hyperspectral delivery check complete:

  • Logsheet is missing
  • proj_tidy.sh complains dem is incorrectly named missing the word ASTER. If this is changed the Readme needs to be updated as well.
  • Mention dropped scan in lines h11, h5, e11 and e5, maybe mention the black dots/lines on line e13 created by the sharp turning.

comment:9 Changed 7 years ago by mark1

Corrected above items - recreated mapped data of line e13 using a larger interpolation distance to fill holes at bend.

comment:10 Changed 7 years ago by besm

Updated lidar data quality report in deliveries before sending.

Version 0, edited 7 years ago by besm (next)

comment:11 Changed 7 years ago by besm

Finished tagging rcd and creating delivery. Ready for checking.

comment:12 Changed 7 years ago by knpa

Beginning DC (with edfi)

comment:13 Changed 7 years ago by knpa

Beginning LiDAR processing.

The flight is too far away from the basestation so having to use PPP result that mark1 got earlier.

comment:14 Changed 7 years ago by knpa

mark1 has created sample LiDAR data for 3 lines, for which he found a roll values of 0.0004.

comment:15 Changed 7 years ago by mark1

  • Description modified (diff)

Camera, Hyperspectral and sample lidar data have been sent to PI at GEUS 16/11/2012.

comment:16 Changed 7 years ago by mark1

SCTs:

line eagle hawk
1 0 -0.03
2 0 0.0
3 0 -0.03
4 0 -0.05
5 0 0.0
6 0 -0.04
7 0 -0.05
8 0 -0.05
9 0 -0.05
10 0 -0.03
11 0 -0.05
12 0 -0.03
13 -0.05 -0.07
14 0 -0.05
15 0 -0.04
16 0 -0.05
17 0 -0.04
18 0 0.70
19 0 0.0
20 0 -0.04
21 0 -0.03

comment:17 Changed 7 years ago by besm

Delivered camera and hyperspectral with a LiDAR sample to Tapsa Tukiainer last Friday (16/11/2012) across disks 134 and 135.

comment:18 Changed 7 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

There are pitch offsets for this flight.

Found a value of -0.00317 for lines 1109 and 1120 (original: -0.00298)

comment:19 Changed 7 years ago by knpa

Found value of -0.00313 for lines 1017 1027

comment:20 Changed 7 years ago by knpa

-0.00317 value fine for 1215 1226

comment:21 Changed 7 years ago by knpa

Original pitch (-0.00298) fine for lines 0934 and 0945

comment:22 Changed 7 years ago by knpa

pitch varies greatly between lines - need to boresight each pair individually :/

comment:23 Changed 7 years ago by knpa

Pitch seems to gradually vary from -0.00298 to -0.00317 throughout flight.

The following values give satisfactory results:

0934: -0.00298
0945: -0.00298
0956: -0.00298
1007: -0.00305
1017: -0.00313
1027: -0.00313
1048: -0.00313
1058: -0.00313
1109: -0.00313
1120: -0.00313
1131: -0.00317
1142: -0.00317
1153: -0.00317
1205: -0.00317
1215: -0.00317
1226: -0.00317
1237: -0.00317
1247: -0.00317
1258: -0.00317
1311: -0.00317

comment:24 Changed 7 years ago by knpa

Forgot 1037: -0.00313

comment:25 Changed 7 years ago by knpa

roll 0.0004 seems to apply to the rest of the flightlines.

Found all height discrepancies between flightlines.

comment:26 Changed 7 years ago by knpa

Finished classifying.

comment:27 Changed 7 years ago by knpa

Delivery created. Ready for delivery check.

comment:28 Changed 7 years ago by edfi

LiDAR on disk 146 to be sent with other data for same project

comment:29 Changed 7 years ago by edfi

Delivery Check of LiDAR completed 5/12/12

comment:30 Changed 7 years ago by edfi

  • Description modified (diff)

comment:31 Changed 7 years ago by mark1

  • Description modified (diff)

comment:32 Changed 6 years ago by edfi

LiDAR delivery check completed before dispatch

comment:33 Changed 6 years ago by emca

Email from Erik Vest Sorensen re: rcd data. He was having trouble with orthorectification.
Checked the event file and the kappa values were incorrect (not offest by -90 degrees).

Re-created the eventfile and sent it to him. Will created another delivery with new eventfile and tiffs tagged with corrected value.

comment:34 Changed 5 years ago by emca

Created a new camera delivery containing the corrected eventfile and the photographs tagged with correct kappa value.

comment:35 Changed 4 years ago by tec

Archival
Started

Ready for upload

Last edited 4 years ago by tec (previous) (diff)

comment:36 Changed 4 years ago by dac

Archiving

Archived data available at NEODC.

comment:37 Changed 4 years ago by dac

  • Resolution set to fixed
  • Status changed from assigned to closed
Note: See TracTickets for help on using tickets.